# SUMO: Subspace-Aware Moment-Orthogonalization for Accelerating Memory-Efficient LLM Training

### v1

@@ Abstract
Low-rank gradient-based optimization methods have significantly improved memory efficiency during the training of large language models (LLMs), enabling operations within constrained hardware without sacrificing performance. However, these methods primarily emphasize memory savings, often overlooking potential acceleration in convergence due to their reliance on standard isotropic steepest descent techniques, which can perform suboptimally in the highly anisotropic landscapes typical of deep networks, particularly LLMs. In this paper, we propose SUMO (Subspace-Aware Moment-Orthogonalization), an optimizer that employs exact singular value decomposition (SVD) for moment orthogonalization within a dynamically adapted low-dimensional subspace, enabling norm-inducing steepest descent optimization steps. By explicitly aligning optimization steps with the spectral characteristics of the loss landscape, SUMO effectively mitigates approximation errors associated with commonly used methods like Newton-Schulz orthogonalization approximation. We theoretically establish an upper bound on these approximation errors, proving their dependence on the condition numbers of moments, conditions we analytically demonstrate are encountered during LLM training. Furthermore, we both theoretically and empirically illustrate that exact orthogonalization via SVD substantially improves convergence rates while reducing overall complexity. Empirical evaluations confirm that SUMO accelerates convergence, enhances stability, improves performance, and reduces memory requirements by up to 20% compared to state-of-the-art methods.
@@ 1 Introduction
Low-rank gradient-based optimization methods have become powerful tools for reducing memory consumption during the pre-training and fine-tuning of large language models (LLMs), often without sacrificing performance, and sometimes even improving it. For instance, while pre-training LLaMA 7B typically requires around 58GB of memory, far exceeding the 24GB available on consumer GPUs like RTX 4090, recent advances, such as those discussed in [1, 2, 3], have demonstrated that LLaMA 7B can now be trained from scratch on a single 24GB GPU without the need for costly memory offloading. Theoretical analysis in [1] attributes this efficiency to the inherent low-rank structure of gradients, which allows for optimization within a significantly reduced latent space. Furthermore, [2] found a consistent decrease in gradient rank throughout training, suggesting that low-rank optimization not only reduces memory usage but also converges toward increasingly compact subspaces.
However, despite these advancements, existing methods mainly focus on memory savings and often overlook the potential for accelerating convergence. Current approaches typically rely on standard steepest descent techniques and operate under implicit assumptions of isotropic geometry, which can hinder efficiency in ill-conditioned settings. This observation motivates our primary objective: to develop a subspace-aware optimizer that leverages low-rank structure while adapting to the loss landscape’s geometry. By reevaluating the choice of norm and its influence on gradient descent dynamics, we aim to design an algorithm that improves generalization, accelerates convergence, while preserving the memory advantages of low-rank methods.
Classical gradient descent, including SGD [4], performs steepest descent under the Euclidean norm, which reflects isotropic curvature. However, deep networks exhibit highly anisotropic loss landscapes, making this assumption suboptimal. Recent work shows that adaptive optimizers like Shampoo [5], SOAP [6], and Muon [7] can be interpreted as steepest descent under non-Euclidean norms tailored to network architecture and data structure. As shown in [8], these methods implicitly adapt to spectral or operator norms, which better capture local curvature and improve convergence. This motivates the design of subspace-aware optimizers that exploit both low-rank structure and appropriate geometry to accelerate training.
[…]
@@
[…]
@@
While norm-induced optimization methods offer a principled way to align updates with the geometry of the loss landscape, their practical deployment often incurs substantial computational overhead. For instance, Shampoo requires computing matrix inverses or root operations at every optimization step, which can be computationally expensive for large-scale neural networks. Similarly, Muon’s first-order moments-orthogonalization, though effective, involves an expensive approximation of spectral decompositions, which is computed by applying five iterations of Newton-Schulz [12] (referred to as Newton-Schulz5). Therefore, there is an inherent trade-off between the theoretical optimality provided by these norm-induced optimization approaches and their practical computational demands.
@@
To bridge the gap between the geometric advantages of norm-induced methods and their computational costs, we first analyze the limitations of existing approximations. We derive an upper bound on the error introduced by the Newton-Schulz orthogonalization, demonstrating that this error increases with the condition number of the moment matrix. This finding explains the increasing instability of the Newton–Schulz5 method in ill-conditioned scenarios, which we subsequently demonstrate to occur in the first-order moment matrices during the training of large language models (LLMs). Building on this analysis, we establish a convergence rate for Muon optimization and compare it to an alternative method that replaces the Newton–Schulz approach with exact Singular Value Decomposition (SVD). Remarkably, we find that the SVD-based approach achieves faster convergence, with improvements directly proportional to the accumulated errors from the moments orthogonalization by the Newton–Schulz5 method. Motivated by the empirical observation that gradients in LLMs often exhibit a low-rank structure, especially during early training, we propose a subspace-aware optimization scheme. This scheme performs exact SVD-based moment orthogonalization within a low-dimensional adaptive subspace. This approach benefits from the relatively low computational cost associated with SVD calculations for low-rank input matrices and enhances the stability of convergence. Also, our approach attains an even greater reduction in memory usage than all previous low-rank training methods by relying solely on first-order moment, as detailed below in Table 1. We support our method with a theoretical convergence guarantee and validate its empirical benefits through experiments, demonstrating faster training and better model performance compared to existing methods.
@@ 2 Related Work
Low-rank gradients naturally emerge during neural network training, as shown in both theoretical and empirical studies [13, 14, 15]. This structure has been leveraged to reduce memory and computational costs during training [16, 17, 18]. Recent work [2] showed that gradients in reversible layers [19] tend to collapse to rank one over time and used this to adaptively adjust gradient rank in Adam. In this paper, we show that the same low-rank trend appears in the first-order moment, which we exploit to apply exact orthogonalization efficiently—avoiding the accumulation of approximation errors, such as Newton-Schultz, during optimization.
@@
Reducing the memory demands of training large language models (LLMs) has driven extensive algorithmic research. One research direction, initiated by LoRA [20], reduces trainable parameters via low-rank adaptation. Yet, such methods often fall short of fully parameterized models, especially during pre-training. Another direction does not restrict the set of trainable parameters but rather optimizes the training methods, with notable examples including AdaRankGrad, GaLore, Fira, Flora, Adam-mini, GaLore-mini, LDAdam, GoLore, LoQT, and Apollo [2, 1, 21, 22, 23, 24, 25, 3], integrating low-rank gradient projections in optimization. In this work, we reduce memory usage even further by relying solely on a first-order momentum, as shown in Table 1.
@@
Preconditioning the Gradient method is critical in enhancing optimizers’ efficiency and effectiveness. Several notable approaches for using a preconditioner have emerged, including methods based on signed gradients [26, 27, 28, 29], gradient clipping [30], normalization [30, 31], and gradient whitening [32, 33, 34, 7, 35, 36, 37]. Recent studies [7, 38] explored gradient orthogonalization strategies, speeding up training. Orthogonalizing gradients effectively constrains updates to lie on directions of uniform magnitude (spectral radius = 1), preventing updates from exaggerating certain gradient directions over others. This procedure ensures a form of normalization that mitigates potential instabilities from ill-conditioned gradients. Unlike these methods, which apply preconditioning or approximate orthogonalization in the high-dimensional parameter space, our approach performs exact SVD-based orthogonalization within an adaptively selected low-rank subspace, offering improved stability and lower computational overhead.
@@
OSGDM [38] is a recently introduced first-order optimization method designed to speed up neural network training by orthogonalizing gradients before the optimization step. Specifically, for a data batch $\xi^{(t)}$, OSGDM applies SVD to the gradient matrix ${\bf G}_{l}=\nabla_{{\bf W}_{l}}\mathcal{L}\left(\Phi\left(\xi^{(t)};\mathbf{\theta}\right)\right)$ of each neural network layer $l$ to generate an orthonormal gradient approximation ${\bf O}_{l}$. This ensures diversity among learned representations and reduces redundancy. The update rule for OSGDM with momentum term $\gamma$ and learning rate $\eta$ is defined as,
$\mathbf{O}_{l}^{(t)}=\operatorname{orth}\left({\bf G}_{l}\right),\quad{\bf M}^{(t+1)}_{l}\leftarrow\gamma{\bf M}^{(t)}_{l}+\eta\mathbf{O}_{l}^{(t)},\quad{\bf W}^{(t+1)}_{l}\leftarrow{\bf W}^{(t)}_{l}-{\bf M}^{(t+1)}_{l},$
where $\operatorname{orth}({\bf G})=\left({\bf G}{\bf G}^{\top}\right)^{-1/2}{\bf G}$ is the ortogonalization operator, and ${\bf M}_{l}$ is the first order moment of layer $l$. Despite additional computational overhead from SVD, OSGDM empirically achieves faster convergence and improved accuracy compared to common methods such as Adam.
@@ 3.1 Theoretical Motivation: Exact moments orthogonalization leads to significantly faster convergence
Previous work on pre-training and fine-tuning large language models (LLMs) has primarily focused on reducing memory usage for constrained hardware or lowering computational cost (for example in [1, 40, 3]). In this paper, we take a step toward accelerating LLM optimization by showing that applying exact orthogonalization (e.g., via SVD) to the first-order moment offers a practical advantage, even over the most accurate approximations, such as the commonly used Newton-Schulz5 method. Specifically, we find that SVD yields faster convergence and lower computational overhead. To support this, we first present a new observation: the moment matrix in LLM training tends to decrease in rank over time. Building on this, we then derive an upper bound on the approximation error of Newton-Schulz5, showing that it depends on both the number of iterations and the matrix condition number, highlighting its limitations in ill-conditioned or low-rank settings (which is exactly the case in LLM optimization moments). This motivates the need for more accurate orthogonalization of moment matrices during LLM training. Of course, applying SVD directly to full-sized layers is generally impractical. The surprising result, however, is that when integrated into a low-rank optimization scheme, the use of SVD becomes not only feasible but preferable. We conclude with a convergence analysis of Muon optimization, which, under these conditions, converges significantly more slowly than the SVD-based alternative. To the best of our knowledge, our convergence analysis of Muon optimization is the first to avoid neglecting the error in the Newton-Schultz approximation [41]. The proofs of all lemmas and theorems of this section are relegated to the Appendix A.
@@
According to the lemma, the approximation error grows exponentially with the condition number. Given the low-rank structure of the first-order moments, low-dimensional optimization offers a means to mitigate this error. Specifically, projecting the moment estimates $\hat{{\bf M}}^{(t)}$ onto their dominant (small) $r$-dimensional subspace ensures that the squared moment $\hat{{\bf M}}^{(t)}\hat{{\bf M}}^{(t)\top}$ is constructed using only the top $r$ squared eigenvalues. These dominant components are significantly larger and exclude near-zero values, resulting in a substantially lower condition number compared to that of the full-rank squared moment matrix. This observation motivates the use of the Muon optimizer within a low-rank optimization framework for LLMs, including 2D reversible layers. Such an approach not only preserves the inherent memory efficiency of low-rank methods but also reduces the approximation error in the optimization step, potentially leading to faster convergence and improved performance over full-dimensional training. However, we also empirically observe that the eigenvalues of the moment matrix decay gradually. As shown in Figure 1, even when projecting onto the dominant subspace, the resulting matrix $\hat{{\bf M}}^{(t)}\hat{{\bf M}}^{(t)\top}$, composed of the top $r=16$ squared eigenvalues, can still exhibit a large condition number, thereby introducing non-negligible approximation error.
@@
To comprehend the cumulative error of Newton-Schulz5 orthogonalization at each optimization step, we proceed to derive the convergence rate of the Moun optimization. To that end, we now provide some notations. Consider a neural network denoted as $\Phi(\cdot;\boldsymbol{\theta})$, which consists of $L$ layers and is parameterized by $\boldsymbol{\theta}\triangleq\left[{\bf W}_{1}^{d_{1}\times d_{0}},\ldots,{\bf W}_{L-1}^{d_{L-1}\times d_{L-2}},{\bf W}_{L}^{d_{L}\times d_{0}^{L-1}}\right]$. Here, ${\bf W}_{i}$ represents the weights tensor parameters associated with the $i$-th layer, for $i\in[L]$. We denote the differential loss $\mathcal{L}$, where with a slight abuse of notation, we write the training problem by $\min_{\mathbf{W}}\ \mathcal{L}(\mathbf{W})=\mathbb{E}_{\xi}[\mathcal{L}(\Phi(\mathbf{W},\xi))],$ if the context refers to the weights of a certain layer. We use the Frobenius norm, denoted $\|\cdot\|_{F}$, which is induced by the inner product $\langle{\bf X},{\bf Y}\rangle=\operatorname{tr}({\bf X}^{\top}{\bf Y})$. Assume that the stochastic gradient $\nabla\mathcal{L}(\mathbf{W},\xi)$ is an unbiased estimator of the full gradient $\nabla\mathcal{L}(\mathbf{W})$, with variance bounded by $\sigma^{2}$, i.e., $\mathbb{E}[\|\nabla\mathcal{L}(\mathbf{W},\xi)-\nabla\mathcal{L}(\mathbf{W})\|_{F}^{2}]\leq\sigma^{2}.$ Let $\mathbf{\mathcal{E}}^{(t)}_{i}=\text{orth}({\bf M}^{(t)})-\text{Newton-Schulz}({\bf M}^{(t)})$ denote the approximation error of the Newton-Schulz (with $i\geq 1$ iteration) at time $t$, where ${\bf M}^{(t)}$ denotes the moment at iteration $t$.
@@
$\mathbf{M}^{(t)}\leftarrow\beta\mathbf{M}^{(t-1)}+(1-\beta)\mathbf{G}_{t}$
[…]
@@
  • There exists $\delta>0$ such that $\|\mathbf{\mathcal{E}}^{(t)}_{i}\|\leq\delta\|{\bf U}_{t}{\bf V}_{t}^{\top}\|=\delta\sqrt{m}$, for all $t$.
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\|\nabla\mathcal{L}(\mathbf{W}^{(t)})\|\right]\leq\mathcal{O}\left(\frac{\sqrt{mRL}}{\sqrt{T}}+\frac{\sigma}{T^{3/2}}+\frac{\sigma}{\sqrt{T}}\right).$
@@
A reduction in $\delta$ is associated with an improvement in the convergence rate. Furthermore, it should be noted that $\delta$ influences the step size $\eta$; a larger $\delta$ results in a smaller step size, providing an additional explanation for the convergence rate.
@@
We acknowledge that the findings of our analysis are applicable only under the conditions specified in $1-4\sqrt{m}\delta>0\Rightarrow\delta<\frac{1}{4\sqrt{m}}$. In scenarios where $\delta>\frac{1}{4\sqrt{m}}$ applies, the algorithm may fail to converge. To ensure that $\delta$ remains sufficiently small, the Newton-Schulz5 method necessitates a substantial number of iterations, consequently slowing down the convergence.

### v2

@@ Abstract
Low-rank gradient-based optimization methods have significantly improved memory efficiency during the training of large language models (LLMs), enabling operations within constrained hardware without sacrificing performance. However, these methods primarily emphasize memory savings, often overlooking potential acceleration in convergence due to their reliance on standard isotropic steepest descent techniques, which can perform suboptimally in the highly anisotropic landscapes typical of deep networks, particularly LLMs. In this paper, we propose SUMO (Subspace-Aware Moment-Orthogonalization), an optimizer that employs exact singular value decomposition (SVD) for moment orthogonalization within a dynamically adapted low-dimensional subspace, enabling norm-inducing steepest descent optimization steps. By explicitly aligning optimization steps with the spectral characteristics of the loss landscape, SUMO effectively mitigates approximation errors associated with commonly used methods, such as the Newton-Schulz orthogonalization approximation. We theoretically establish an upper bound on these approximation errors, proving their dependence on the condition numbers of moments, conditions we analytically demonstrate are encountered during LLM training. Furthermore, we both theoretically and empirically illustrate that exact orthogonalization via SVD substantially improves convergence rates while reducing overall complexity. Empirical evaluations confirm that SUMO accelerates convergence, enhances stability, improves performance, and reduces memory requirements by up to 20% compared to state-of-the-art methods.
@@ 1 Introduction
Low-rank gradient-based optimization methods have become powerful tools for reducing memory consumption during the pre-training and fine-tuning of large language models (LLMs), often without sacrificing performance, and sometimes even improving it. For instance, while pre-training LLaMA 7B typically requires around 58GB of memory, far exceeding the 24GB available on consumer GPUs like RTX 4090, recent advances, such as those discussed in [1, 2, 3], have demonstrated that LLaMA 7B can now be trained from scratch on a single 24GB GPU without the need for costly memory offloading. The theoretical analysis in [1] attributes this efficiency to the inherent low-rank structure of the gradients, which enables optimization in a significantly reduced latent space. Furthermore, [2] found a consistent decrease in gradient rank throughout training, suggesting that low-rank optimization not only reduces memory usage but also converges toward increasingly compact subspaces.
However, despite these advancements, existing methods primarily focus on memory savings and often overlook the potential to accelerate convergence. Current approaches typically rely on standard steepest descent methods and assume isotropic geometry, which can hinder efficiency in ill-conditioned settings. This observation motivates our primary objective: to develop a subspace-aware optimizer that leverages low-rank structure while adapting to the geometry of the loss landscape. By reevaluating the choice of norm and its influence on gradient descent dynamics, we aim to design an algorithm that improves generalization, accelerates convergence, while preserving the memory advantages of low-rank methods.
Classical gradient descent, including SGD [4], performs steepest descent under the Euclidean norm, which reflects isotropic curvature. However, deep networks exhibit highly anisotropic loss landscapes, making this assumption suboptimal. Recent work shows that adaptive optimizers, such as Shampoo [5], SOAP [6], and Muon [7], can be interpreted as steepest descent under non-Euclidean norms tailored to the network architecture and data structure. As shown in [8], these methods implicitly adapt to spectral or operator norms, which better capture local curvature and improve convergence. This motivates the design of subspace-aware optimizers that exploit both low-rank structure and appropriate geometry to accelerate training.
To formalize the role of geometry in optimization, consider a neural network with a differentiable loss function $\mathcal{L}:\mathcal{W}\to\mathbb{R}$ defined on a weight space $\mathcal{W}=\mathbb{R}^{n}$. The local behavior around a point $\mathbf{w}$ can be approximated by the Taylor expansion, $\mathcal{L}(\mathbf{w}+\Delta\mathbf{w})\approx\mathcal{L}(\mathbf{w})+\mathbf{g}^{\top}\Delta\mathbf{w}+\frac{\lambda}{2}\|\Delta\mathbf{w}\|^{2},$ where $\mathbf{g}=\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{w})$, $\lambda>0$ captures the sharpness or curvature of the loss surface and $\|\cdot\|$ is a chosen norm reflecting the geometry of the optimization landscape. Minimizing this approximation corresponds precisely to performing steepest descent under the given norm constraint. According to [9], the solution to this minimization explicitly takes the form,
@@
where $\|\cdot\|_{*}$ denotes the dual norm of $\|\cdot\|$. Adaptive optimizers differ primarily in their norm choices. Adam utilizes a dynamic Max-of-Max norm constraint. Recent optimizers consider matrix norms while applying steepest descent at the layer level. Muon imposes a fixed Schatten-$p$ norm constraint for large $p$, effectively using the spectral norm on weight matrices [10, 7]. Shampoo [5] dynamically learns the optimal approximate Schatten-$p$ norm for steepest descent, with its variants like SOAP [11] apply momentum to efficiently navigate the space of possible norms. Muon, by contrast, operates within a relatively fixed but large Schatten-$p$ norm, striking a balance between the dynamic adaptability of Shampoo and the static spectral norm constraints. Since neural network weights locally act as linear operators on Euclidean spaces, the induced operator (spectral) norm provides a natural constraint that aligns with the curvature of the loss surface. This perspective motivates gradient orthogonalization, which ensures that optimization updates respect the spectral norm, thereby controlling perturbation magnitude and enhancing optimization stability and efficiency [9].
@@
While norm-induced optimization methods offer a principled way to align updates with the geometry of the loss landscape, their practical deployment often incurs substantial computational overhead. For instance, Shampoo requires computing matrix inverses or root operations at every optimization step, which can be computationally expensive for large-scale neural networks. Similarly, Muon’s first-order moments-orthogonalization, though effective, involves a costly approximation to spectral decompositions, computed by applying five iterations of Newton-Schulz [12] (referred to as Newton-Schulz5). Therefore, there is an inherent trade-off between the theoretical optimality provided by these norm-induced optimization approaches and their practical computational demands.
@@
To bridge the gap between the geometric advantages of norm-induced methods and their computational costs, we first analyze the limitations of existing approximations. We derive an upper bound on the error introduced by the Newton-Schulz orthogonalization, demonstrating that this error increases with the condition number of the moment matrix. This finding explains the increasing instability of the Newton–Schulz5 method in ill-conditioned scenarios, which we subsequently demonstrate to occur in the first-order moment matrices during the training of large language models (LLMs). Building on this analysis, we establish a convergence rate for Muon optimization and compare it to an alternative method that replaces the Newton–Schulz approach with exact Singular Value Decomposition (SVD). Remarkably, we find that the SVD-based approach converges faster, with improvements directly proportional to the accumulated errors from the moments’ orthogonalization by the Newton–Schulz5 method. Motivated by the empirical observation that gradients in LLMs often exhibit a low-rank structure, especially during early training, we propose a subspace-aware optimization scheme. This scheme performs exact SVD-based moment orthogonalization within a low-dimensional adaptive subspace. This approach benefits from the relatively low computational cost of SVD calculations for low-rank input matrices and enhances convergence stability. Also, our approach attains an even greater reduction in memory usage than all previous low-rank training methods by relying solely on the first-order moment, as detailed below in Table 1. We support our method with a theoretical convergence guarantee and validate its empirical benefits through experiments, demonstrating faster training and better model performance compared to existing methods.
@@ 2 Related Work
Low-rank gradients naturally emerge during neural network training, as shown in both theoretical and empirical studies [13, 14, 15]. This structure has been leveraged to reduce memory and computational costs during training [16, 17, 18]. Recent work [2] showed that gradients in reversible layers [19] tend to collapse to rank one over time and used this to adaptively adjust gradient rank in Adam. In this paper, we demonstrate that the same low-rank trend is present in the first-order moment, which we utilize to efficiently apply exact orthogonalization—avoiding the accumulation of approximation errors, such as those encountered in Newton-Schultz, during optimization.
@@
Reducing the memory demands of training large language models (LLMs) has driven extensive algorithmic research. One research direction, initiated by LoRA [20], reduces the number of trainable parameters via low-rank adaptation. Yet, such methods often fall short of fully parameterized models, especially during pre-training. Another direction does not restrict the set of trainable parameters but instead optimizes the training methods, with notable examples including AdaRankGrad, GaLore, Fira, Flora, Adam-mini, GaLore-mini, LDAdam, GoLore, LoQT, and Apollo [1, 2, 3, 21, 22, 23, 24, 25], integrating low-rank gradient projections in optimization. In this work, we reduce memory usage even further by relying solely on a first-order momentum, as shown in Table 1.
@@
Preconditioning the Gradient method is critical in enhancing the efficiency and effectiveness of optimizers. Several notable approaches for using a preconditioner have emerged, including methods based on signed gradients [26, 27, 28, 29], gradient clipping [30], normalization [30, 31], and gradient whitening [7, 32, 33, 34, 35, 36, 37]. Recent studies [7, 38] explored gradient-orthogonalization strategies, thereby speeding up training. Orthogonalizing gradients effectively constrains updates to lie on directions of uniform magnitude (spectral radius = 1), preventing updates from exaggerating specific gradient directions over others. This procedure ensures a form of normalization that mitigates potential instabilities from ill-conditioned gradients. Unlike these methods, which apply preconditioning or approximate orthogonalization in the high-dimensional parameter space, our approach performs exact SVD-based orthogonalization within an adaptively selected low-rank subspace, offering improved stability and lower computational overhead.
@@
OSGDM [38] is a recently introduced first-order optimization method that speeds up neural network training by orthogonalizing gradients before the optimization step. Specifically, for a data batch $\xi^{(t)}$, OSGDM applies SVD to the gradient matrix ${\bf G}_{l}^{(t)}=\nabla_{{\bf W}_{l}}\mathcal{L}\left(\Phi\left(\xi^{(t)};\mathbf{\theta}\right)\right)$ of each neural network layer $l$ to generate an orthonormal gradient approximation ${\bf O}_{l}$. This ensures diversity among learned representations and reduces redundancy. The update rule for OSGDM with momentum term $\gamma$ and learning rate $\eta$ is defined as,
$\mathbf{O}_{l}^{(t)}=\operatorname{orth}\left({\bf G}_{l}^{(t)}\right),\quad{\bf M}^{(t+1)}_{l}\leftarrow\gamma{\bf M}^{(t)}_{l}+\eta\mathbf{O}_{l}^{(t)},\quad{\bf W}^{(t+1)}_{l}\leftarrow{\bf W}^{(t)}_{l}-{\bf M}^{(t+1)}_{l}$,
where $\operatorname{orth}({\bf G})=\left({\bf G}{\bf G}^{\top}\right)^{-1/2}{\bf G}$ is the ortogonalization operator, and ${\bf M}_{l}$ is the first order moment of layer $l$. Despite the additional computational overhead of SVD, OSGDM empirically converges faster and achieves higher accuracy than common methods such as Adam.
@@ 3.1 Theoretical Motivation: Exact moments orthogonalization leads to significantly faster convergence
Previous work on pre-training and fine-tuning large language models (LLMs) has primarily focused on reducing memory usage for constrained hardware or lowering computational cost (e.g., [1, 40, 3]). In this paper, we take a step toward accelerating LLM optimization by showing that applying exact orthogonalization (e.g., via SVD) to the first-order moment offers a practical advantage, even over the most accurate approximations, such as the commonly used Newton-Schulz5 method. Specifically, we find that SVD converges faster and incurs lower computational overhead. To support this, we first present a new observation: the moment matrix in LLM training tends to decrease in rank over time. Building on this, we then derive an upper bound on the approximation error of Newton-Schulz5, showing that it depends on both the number of iterations and the matrix condition number, highlighting its limitations in ill-conditioned or low-rank settings (which is precisely the case in LLM optimization moments). This motivates the need for more accurate orthogonalization of moment matrices during LLM training. Of course, applying SVD directly to full-sized layers is generally impractical. The surprising result, however, is that when integrated into a low-rank optimization scheme, the use of SVD becomes not only feasible but preferable. We conclude with a convergence analysis of Muon optimization, which, under these conditions, converges significantly more slowly than the SVD-based alternative. To the best of our knowledge, our convergence analysis of Muon optimization is the first to avoid neglecting the error in the Newton-Schultz approximation [41]. The proofs of all lemmas and theorems of this section are relegated to the Appendix A.
@@
According to the lemma, the approximation error grows exponentially with the condition number. Given the low-rank structure of the first-order moments, low-dimensional optimization can mitigate this error. Specifically, projecting the moment estimates $\hat{{\bf M}}^{(t)}$ onto their dominant (small) $r$-dimensional subspace ensures that the squared moment $\hat{{\bf M}}^{(t)}\hat{{\bf M}}^{(t)\top}$ is constructed using only the top $r$ squared eigenvalues. These dominant components are significantly larger and exclude near-zero values, resulting in a substantially lower condition number compared to that of the full-rank squared moment matrix. This observation motivates the use of the Muon optimizer within a low-rank optimization framework for LLMs, including 2D reversible layers. Such an approach not only preserves the inherent memory efficiency of low-rank methods but also reduces approximation error during optimization, potentially leading to faster convergence and improved performance compared to full-dimensional training. However, we also empirically observe that the eigenvalues of the moment matrix decay gradually. As shown in Figure 1, even when projecting onto the dominant subspace, the resulting matrix $\hat{{\bf M}}^{(t)}\hat{{\bf M}}^{(t)\top}$, composed of the top $r=16$ squared eigenvalues, can still exhibit a large condition number, thereby introducing non-negligible approximation error.
@@
To comprehend the cumulative error of Newton-Schulz5 orthogonalization at each optimization step, we proceed to derive the convergence rate of the Moun optimization. To that end, we now provide some notations. Consider a neural network denoted as $\Phi(\cdot;\boldsymbol{\theta})$, which consists of $L$ layers and is parameterized by $\boldsymbol{\theta}\triangleq\left[{\bf W}_{1}^{d_{1}\times d_{0}},\ldots,{\bf W}_{L-1}^{d_{L-1}\times d_{L-2}},{\bf W}_{L}^{d_{L}\times d_{L-1}}\right]$. Here, ${\bf W}_{i}$ represents the weights tensor parameters associated with the $i$-th layer, for $i\in[L]$. We denote the differential loss $\mathcal{L}$, where, with a slight abuse of notation, we write the training problem by $\min_{\mathbf{W}}\ \mathcal{L}(\mathbf{W})=\mathbb{E}_{\xi}[\mathcal{L}(\Phi(\mathbf{W},\xi))],$ if the context refers to the weights of a certain layer. We use the Frobenius norm, denoted $\|\cdot\|_{F}$, which is induced by the inner product $\langle{\bf X},{\bf Y}\rangle=\operatorname{tr}({\bf X}^{\top}{\bf Y})$. Assume that the stochastic gradient $\nabla\mathcal{L}(\mathbf{W},\xi)$ is an unbiased estimator of the full gradient $\nabla\mathcal{L}(\mathbf{W})$, with variance bounded by $\sigma^{2}$, i.e., $\mathbb{E}[\|\nabla\mathcal{L}(\mathbf{W},\xi)-\nabla\mathcal{L}(\mathbf{W})\|_{F}^{2}]\leq\sigma^{2}.$ Let $\mathbf{\mathcal{E}}^{(t)}_{i}=\text{orth}({\bf M}^{(t)})-\text{Newton-Schulz}({\bf M}^{(t)})$ denote the approximation error of the Newton-Schulz (with $i\geq 1$ iteration) at time $t$, where ${\bf M}^{(t)}$ denotes the moment at iteration $t$.
@@
$\mathbf{M}^{(t)}\leftarrow\beta\mathbf{M}^{(t-1)}+(1-\beta)\mathbf{G}^{(t)}$
$\mathbf{O}^{(t)}\leftarrow\mathbf{U}^{(t)}\mathbf{V}^{(t)^{\top}}+\mathcal{E}_{i}^{(t)},\quad(i\text{ iterations Newton-Schulz approximation })$
@@
  • There exists $\delta>0$ such that $\|\mathbf{\mathcal{E}}^{(t)}_{i}\|\leq\delta\|\mathbf{U}^{(t)}{\mathbf{V}^{(t)}}^{\top}\|=\delta\sqrt{n}$, for all $t$.
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\|\nabla\mathcal{L}(\mathbf{W}^{(t)})\|\right]\leq\mathcal{O}\left(\frac{\sqrt{nRL}}{\sqrt{T}}+\frac{\sigma}{T^{3/2}}+\frac{\sigma}{\sqrt{T}}\right).$
@@
A reduction in $\delta$ is associated with an improvement in the convergence rate. Furthermore, it should be noted that $\delta$ influences the step size $\eta$; a larger $\delta$ yields a smaller $\eta$, providing an additional explanation for the convergence rate.
@@
We acknowledge that the findings of our analysis are applicable only under the conditions specified in $1-4\sqrt{n}\delta>0\Rightarrow\delta<\frac{1}{4\sqrt{n}}$. In scenarios where $\delta>\frac{1}{4\sqrt{n}}$ applies, the algorithm may fail to converge. To ensure that $\delta$ remains sufficiently small, the Newton-Schulz5 method necessitates a substantial number of iterations, consequently slowing down the convergence.